# SCoTER: Structured Chain-of-Thought Transfer for Enhanced Recommendation

### v1

@@ 1. Introduction
Large Language Models (LLMs) (llmsurvey1; llmsurvey2) have demonstrated strong reasoning capabilities, especially with Chain-of-Thought (CoT) prompting (wei2022chain). However, transferring this power from objective, logic-driven tasks to the subjective domain of recommender systems introduces two fundamental and interdependent challenges. The first challenge is determining what to transfer. Defining and discovering effective reasoning patterns is inherently difficult without a clear ground truth. User intents are highly diverse, meaning no single pattern can generalize, and sparse data for long-tail items further complicates the validation of any proposed reasoning path. This reveals an inherent conflict between two objectives: preserving the stepwise CoT structure, the source of its reasoning power, and satisfying the stringent low-latency demands of production environments. Addressing these intertwined problems requires a framework that jointly addresses reasoning discovery and structure-preserving transfer.
@@
Despite the clear need for a unified framework, no such solution currently exists. Current approaches address these challenges in isolation, which can be doubly flawed. First, the standalone solutions are inherently inadequate. Pattern discovery for the ’what to transfer’ problem has largely been guided by heuristic, handcrafted schemas conceived from prior assumptions rather than empirical results. This practice decouples pattern design from the downstream recommendation task, resulting in brittle patterns that fail to generalize (wang2023self; zhangautomatic; wang2023plan). Meanwhile, transfer mechanisms for the ’how to transfer’ problem have typically sacrificed the structural integrity of reasoning, treating it as static features divorced from generative logic (wang2024can; xi2024towards; yang2025reccot; xia2025trackrec). Second, and more critically, this very separation prevents any possibility of joint optimization. It fosters a problematic cycle where patterns are designed without regard for their integration costs, while integration strategies are developed without preserving the patterns’ core logical effectiveness.
@@ 2. Related works
LLM Reasoning for Recommendation. Recent approaches integrating LLM into recommendation systems have explored a variety of complex reasoning structures. CoT-Rec (liu2025improving) employs two-stage prompting for user preference analysis, GOT4Rec (long2024got4rec) uses Graph-of-Thought frameworks, and ThinkRec (yu2025thinkrec) shifts to System 2 thinking through reasoning data synthesis, while RecGPT (yi2025recgpt) works to unify multi-step reasoning frameworks. Complementing these efforts, a parallel line of research focuses on refining or distilling reasoning capabilities. This includes the inference-time autoregressive refinement in ReaRec (rearec), the distillation of step-by-step rationales to smaller models by RDRec (rdrec), and the iterative feedback framework used by TrackRec (xia2025trackrec). However, these methods are limited by relying on heuristic reasoning paths instead of mining user sequences, and their failure to jointly optimize pattern discovery and integration.
Automated Reasoning Discovery: Automated discovery of reasoning patterns has emerged as an alternative to manual template design. Auto-CoT (zhangautomatic) automatically constructs demonstrations by sampling diverse questions and generating rationales, while Self-prompted CoT (wang2023self) enables LLMs to self-induce reasoning steps. Self-Consistency (wangself) improves reasoning by sampling multiple paths, and broader approaches include APE (zhou2022large) for automatic prompt engineering, PromptBreeder (fernando2024promptbreeder) for evolutionary optimization, and Self-discover (zhou2024self) for composing atomic reasoning modules. However, these methods are primarily designed for objective tasks with verifiable ground truth. They are less effective in the recommendation domain, which is subjective and has sparse rewards that make it difficult to improve reasoning paths. Our approach, in contrast, addresses this by sampling from broad user behaviors and performing in-depth analysis to use the Recall metric as a dense reward signal.
@@ 4.1. What to Transfer: Automated Discovery of Reasoning patterns
Generate: The Generate phase produces a diverse set of candidate reasoning chains for each user sequence, S. We employ an LLM, such as DeepSeek-R1 (guo2025deepseek), with a structured prompt (Figure 4) that instructs the model to act as a ”recommendation expert” and defines a specific output format.
@@
The mining process begins by transforming textual reasoning chains into a dense embedding space using a pre-trained sentence encoder (e.g., Qwen3-8B-Embedding (zhang2025qwen3)). Within this space, we perform unsupervised clustering to group semantically similar chains, forming a set of initial candidate patterns.
@@ 4.2.1. Structured Distillation
A smaller, more efficient student model (Qwen3-8B (yang2025qwen3)) is fine-tuned on this synthetic dataset, enabling it to generate pattern-consistent reasoning chains that adapt to specific user contexts.
@@
For each generated reasoning step $c_{i,j}$, we extract a dense embedding using a pre-trained sentence encoder (e.g., Qwen3-8B-Embedding (zhang2025qwen3)). This process transforms textual reasoning steps into fixed-dimensional embeddings, where $\mathbf{e}_{i,j}\in\mathbb{R}^{D}$ represents the embedding for the $j$-th reasoning step of sequence $S_{i}$:
@@ 4.2.2. Order-Preserving Fusion
To align the reasoning space with the recommendation objective, we employ a contrastive learning component using the InfoNCE loss (oord2018representation). The loss is computed between the final reasoning step embedding, $\mathbf{z}_{K}$, and the target item embedding, $\mathbf{v}_{\text{target}}$:
@@ 5.1. Experimental Setup
Datasets. We conduct experiments on four widely used datasets: three subsets of the Amazon Product Reviews dataset (he2016ups; ni2019justifying) (Beauty, Instruments, and Sports) and the Yelp dataset. Table 2 summarizes the statistics of these datasets. Following previous work (zhou2020s3), we process the data to enforce a 5-core density, removing all users and items with fewer than five interactions. All user sequences are then normalized to a uniform length of 20 through padding or truncation, preserving their most recent interactions. For evaluation, we use the leave-one-out protocol: each user’s final interaction is designated for testing, the penultimate one for validation, and the remaining interactions are used for training
@@
  • MF (koren2009matrix): A classic model that uses matrix factorization to learn latent embeddings for users and items.
  • LightGCN (he2020lightgcn): A graph convolutional network that captures collaborative signals via neighborhood aggregation.
  • Caser (tang2018personalized): A sequential model that employs convolutional neural networks to capture local sequential patterns.
  • HGN (ma2019hierarchical): A sequential model that utilizes a hierarchical gating network to adaptively integrate a user’s long- and short-term preferences.
  • SASRec (kang2018self): A sequential model that uses a self-attention mechanism to capture long-range dependencies and dynamic user preferences.
  • Bert4Rec (sun2019bert4rec): A sequential model that uses a deep bidirectional self-attention mechanism to model user sequences.
  • TIGER (rajput2023recommender): A generative model that represents items as discrete token sequences, enabling recommendation through autoregressive decoding. We select TIGER as the backbone for our method due to its strong generative performance and architectural compatibility with reasoning integration.
@@
Implementation Details. For traditional methods, we follow standard implementations with hyperparameters tuned on validation sets. For generative methods, we adopted a unified configuration based on the T5 architecture. The backbone is a 4-layer Transformer, configured with a model dimension of 128, six attention heads (dimension 64), a 1024-unit hidden MLP, ReLU activation, and a 0.1 dropout rate. The tokenizer employs RQ-VAE for discrete semantic encoding with 4 codebooks, each containing 256 embeddings of dimension 32. Semantic inputs to RQ-VAE are derived from the embeddings of item titles and descriptions processed by Qwen3-8B-Embedding (zhang2025qwen3). During inference, we use a beam size of 20 to balance recommendation quality and efficiency.
@@ 5.6. Online A/B Test
We refer interested readers to (zhang2025gpr) for a more comprehensive discussion of the online experimental setup, additional metrics, and extended analysis.

### v3

@@ 1. Introduction
Large Language Models (LLMs) (Zhao et al., 2023; Huang and Chang, 2023) have demonstrated strong reasoning capabilities, especially with Chain-of-Thought (CoT) prompting (Wei et al., 2022). However, transferring this power from objective, logic-driven tasks to the subjective domain of recommender systems introduces two fundamental and interdependent challenges. The first challenge is determining what to transfer. Defining and discovering effective reasoning patterns is inherently difficult without a clear ground truth. User intents are highly diverse, meaning no single pattern can generalize, and sparse data for long-tail items further complicates the validation of any proposed reasoning path. This reveals an inherent conflict between two objectives: preserving the stepwise CoT structure, the source of its reasoning power, and satisfying the stringent low-latency demands of production environments. Addressing these intertwined problems requires a framework that jointly addresses reasoning discovery and structure-preserving transfer.
@@
Despite the clear need for a unified framework, no such solution currently exists. Current approaches address these challenges in isolation, which can be doubly flawed. First, the standalone solutions are inherently inadequate. Pattern discovery for the ’what to transfer’ problem has largely been guided by heuristic, handcrafted schemas conceived from prior assumptions rather than empirical results. This practice decouples pattern design from the downstream recommendation task, resulting in brittle patterns that fail to generalize (Wang et al., 2023a; Zhang et al., ; Wang et al., 2023b). Meanwhile, transfer mechanisms for the ’how to transfer’ problem have typically sacrificed the structural integrity of reasoning, treating it as static features divorced from generative logic (Wang et al., 2024b; Xi et al., 2024; Yang et al., 2025b; Xia et al., 2025). Second, and more critically, this very separation prevents any possibility of joint optimization. It fosters a problematic cycle where patterns are designed without regard for their integration costs, while integration strategies are developed without preserving the patterns’ core logical effectiveness.
@@ 2. Related works
LLM Reasoning for Recommendation. Recent approaches integrating LLM into recommendation systems have explored a variety of complex reasoning structures. CoT-Rec (Liu et al., 2025) employs two-stage prompting for user preference analysis, GOT4Rec (Long et al., 2024) uses Graph-of-Thought frameworks, and ThinkRec (Yu et al., 2025) shifts to System 2 thinking through reasoning data synthesis, while RecGPT (Yi et al., 2025) works to unify multi-step reasoning frameworks. Complementing these efforts, a parallel line of research focuses on refining or distilling reasoning capabilities. This includes the inference-time autoregressive refinement in ReaRec (Tang et al., 2025), the distillation of step-by-step rationales to smaller models by RDRec (Wang et al., 2024a), and the iterative feedback framework used by TrackRec (Xia et al., 2025). However, these methods are limited by relying on heuristic reasoning paths instead of mining user sequences, and their failure to jointly optimize pattern discovery and integration.
Automated Reasoning Discovery: Automated discovery of reasoning patterns has emerged as an alternative to manual template design. Auto-CoT (Zhang et al., ) automatically constructs demonstrations by sampling diverse questions and generating rationales, while Self-prompted CoT (Wang et al., 2023a) enables LLMs to self-induce reasoning steps. Self-Consistency (Wang et al., ) improves reasoning by sampling multiple paths, and broader approaches include APE (Zhou et al., 2022) for automatic prompt engineering, PromptBreeder (Fernando et al., 2024) for evolutionary optimization, and Self-discover (Zhou et al., 2024) for composing atomic reasoning modules. However, these methods are primarily designed for objective tasks with verifiable ground truth. They are less effective in the recommendation domain, which is subjective and has sparse rewards that make it difficult to improve reasoning paths. Our approach, in contrast, addresses this by sampling from broad user behaviors and performing in-depth analysis to use the Recall metric as a dense reward signal.
@@ 4.1. What to Transfer: Automated Discovery of Reasoning patterns
Generate: The Generate phase produces a diverse set of candidate reasoning chains for each user sequence, S. We employ an LLM, such as DeepSeek-R1 (Guo et al., 2025), with a structured prompt (Figure 4) that instructs the model to act as a ”recommendation expert” and defines a specific output format.
@@
The mining process begins by transforming textual reasoning chains into a dense embedding space using a pre-trained sentence encoder (e.g., Qwen3-8B-Embedding (Zhang et al., 2025b)). Within this space, we perform unsupervised clustering to group semantically similar chains, forming a set of initial candidate patterns.
@@ 4.2.1. Structured Distillation
A smaller, more efficient student model (Qwen3-8B (Yang et al., 2025a)) is fine-tuned on this synthetic dataset, enabling it to generate pattern-consistent reasoning chains that adapt to specific user contexts.
@@
For each generated reasoning step $c_{i,j}$, we extract a dense embedding using a pre-trained sentence encoder (e.g., Qwen3-8B-Embedding (Zhang et al., 2025b)). This process transforms textual reasoning steps into fixed-dimensional embeddings, where $\mathbf{e}_{i,j}\in\mathbb{R}^{D}$ represents the embedding for the $j$-th reasoning step of sequence $S_{i}$:
@@ 4.2.2. Order-Preserving Fusion
To align the reasoning space with the recommendation objective, we employ a contrastive learning component using the InfoNCE loss (Oord et al., 2018). The loss is computed between the final reasoning step embedding, $\mathbf{z}_{K}$, and the target item embedding, $\mathbf{v}_{\text{target}}$:
@@ 5.1. Experimental Setup
Datasets. We conduct experiments on four widely used datasets: three subsets of the Amazon Product Reviews dataset (He and McAuley, 2016; Ni et al., 2019) (Beauty, Instruments, and Sports) and the Yelp dataset. Table 2 summarizes the statistics of these datasets. Following previous work (Zhou et al., 2020), we process the data to enforce a 5-core density, removing all users and items with fewer than five interactions. All user sequences are then normalized to a uniform length of 20 through padding or truncation, preserving their most recent interactions. For evaluation, we use the leave-one-out protocol: each user’s final interaction is designated for testing, the penultimate one for validation, and the remaining interactions are used for training
@@
  • MF (Koren et al., 2009): A classic model that uses matrix factorization to learn latent embeddings for users and items.
  • LightGCN (He et al., 2020): A graph convolutional network that captures collaborative signals via neighborhood aggregation.
  • Caser (Tang and Wang, 2018): A sequential model that employs convolutional neural networks to capture local sequential patterns.
  • HGN (Ma et al., 2019): A sequential model that utilizes a hierarchical gating network to adaptively integrate a user’s long- and short-term preferences.
  • SASRec (Kang and McAuley, 2018): A sequential model that uses a self-attention mechanism to capture long-range dependencies and dynamic user preferences.
  • Bert4Rec (Sun et al., 2019): A sequential model that uses a deep bidirectional self-attention mechanism to model user sequences.
  • TIGER (Rajput et al., 2023): A generative model that represents items as discrete token sequences, enabling recommendation through autoregressive decoding. We select TIGER as the backbone for our method due to its strong generative performance and architectural compatibility with reasoning integration.
@@
Implementation Details. For traditional methods, we follow standard implementations with hyperparameters tuned on validation sets. For generative methods, we adopted a unified configuration based on the T5 architecture. The backbone is a 4-layer Transformer, configured with a model dimension of 128, six attention heads (dimension 64), a 1024-unit hidden MLP, ReLU activation, and a 0.1 dropout rate. The tokenizer employs RQ-VAE for discrete semantic encoding with 4 codebooks, each containing 256 embeddings of dimension 32. Semantic inputs to RQ-VAE are derived from the embeddings of item titles and descriptions processed by Qwen3-8B-Embedding (Zhang et al., 2025b). During inference, we use a beam size of 20 to balance recommendation quality and efficiency.
@@ 5.6. Online A/B Test
We refer interested readers to (Zhang et al., 2025a) for a more comprehensive discussion of the online experimental setup, additional metrics, and extended analysis.